### v1

# Can a Crow Hatch a Falcon? Lineage Matters in Predicting Large Language Model Performance

Abstract
Accurately forecasting the performance of Large Language Models (LLMs) before extensive fine-tuning or merging can substantially reduce both computational expense and development time. Although prior approaches like scaling laws account for global factors such as parameter size or training tokens, they often overlook explicit lineage relationships—i.e., which models are derived or merged from which parents. In this work, we propose a novel Lineage-Regularized Matrix Factorization (LRMF) framework that encodes ancestral ties among LLMs via a graph Laplacian regularizer. By leveraging multi-hop parent–child connections, LRMF consistently outperforms conventional matrix factorization and collaborative filtering methods in both instance-level and benchmark-level performance prediction. Our large-scale study includes 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks, showing that lineage constraints yield up to 7–10 percentage points higher correlation with actual performance compared to baselines. Moreover, LRMF effectively addresses the cold-start problem, providing accurate estimates for newly derived or merged models even with minimal data. This lineage-guided strategy thus offers a resource-efficient way to inform hyperparameter tuning, data selection, and model combination in modern LLM development.
1 Introduction
Building and refining a Large Language Model (LLM) is an increasingly expensive venture. Developers face critical decisions such as model size, the amount and type of training data, and how to allocate parameters for fine-tuning. Moreover, model merging (or model “souping”) is emerging as a promising strategy: distinct models or checkpoints are combined at the parameter level to create new variants with complementary strengths (Wortsman et al., 2022). Each of these steps—training, fine-tuning, and merging—often demands large-scale exploration and incurs a high computational cost.
If one could accurately predict the performance of a proposed LLM variant—even before running large-scale optimization—it would profoundly reduce the search space. Such foresight would streamline choosing (1) the right model size and data volume for pre-training, (2) fine-tuning hyperparameters and data composition (e.g., domain mix), and (3) whether combining certain model checkpoints is worthwhile. Indeed, prior work has sought to provide partial solutions. Classic scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) predict an LLM’s loss on the training corpus as a function of parameter size and token count. These laws help developers extrapolate the compute needed to reach a desired perplexity or training loss. Subsequent studies extend these ideas to specific benchmarks, e.g., MMLU (Hendrycks et al., 2021a; Wang et al., 2024), BBH (Suzgun et al., 2023), or MATH (Hendrycks et al., 2021b), by fitting regression models or matrix factorization frameworks to empirical performance data (Zhang et al., 2024; Polo et al., 2024; Kipnis et al., 2025).
However, real-world LLMs rarely arise in isolation. In practice, a single base model often spawns multiple variants via continual pre-training, domain-specific fine-tuning, preference tuning (e.g., RLHF), or parameter-level merging of multiple checkpoints. Such derivation or lineage relationships typically imply that “child” models inherit certain performance characteristics from their parents. Despite this, previous performance-prediction work has largely ignored the explicit structure of these relationships, focusing instead on coarse factors such as model size or training tokens. As a result, the potential advantage of modeling multi-hop ancestry—i.e., “who was the parent’s parent?”—remains underexplored.
In this paper, we propose a novel way to incorporate such lineage information into performance prediction. Specifically, we introduce a Lineage-Regularized Matrix Factorization (LRMF) framework that treats the known derivation paths among models as a lineage graph. Within a standard matrix factorization procedure (where models form the rows and benchmarks/tasks form the columns), we add a graph Laplacian regularization term that explicitly constrains models with direct or multi-hop parent–child ties to lie “close” in the latent space. The resulting method can more accurately propagate performance signals from well-evaluated parent models to their lesser-known descendants.
We conduct a large-scale empirical study encompassing 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks. We demonstrate that incorporating lineage constraints—i.e., explicitly modeling which models were fine-tuned or merged from which parents—consistently boosts predictive accuracy. In fact, compared to standard matrix factorization baselines, our method improves correlation with true performance by up to 7–10 percentage points on certain tasks, affirming that lineage is a key factor in understanding and forecasting LLM behaviors.
Moreover, by leveraging explicit lineage relationships (parent–child or merges), our method LRMF tackles the cold-start problem, enabling accurate performance estimates for newly derived or merged models without costly full re-benchmarking. Indeed, our experiments reveal that even with as few as 10 labeled instances per model, we can achieve near state-of-the-art correlation with fully evaluated baselines—a remarkable improvement over standard collaborative filtering or scaling-based approaches. Taken together, our findings suggest that preserving and exploiting the lineage graph is vital for guiding resource-efficient decisions in LLM development.
2 Related Work
Performance Estimation via Scaling Laws.
Early studies on large language models (LLMs) have shown that core metrics like cross-entropy loss exhibit approximate power-law relationships with key design factors such as model size, dataset scale, and overall computational budget (Kaplan et al., 2020; Hoffmann et al., 2022; Hernandez et al., 2022; Gordon et al., 2021). While these classical “scaling laws” provide valuable insights into how performance grows with parameters and tokens, their estimates typically assume a single model family or a narrowly defined training setup. Consequently, scaling relationships learned from one architecture (e.g., Transformer-based language models) may not directly generalize to another.
Recent work expands scaling laws beyond a single family. For instance, Ye et al. (2023) investigate whether past performance records from multiple LLM variants can predict new model settings; Owen (2024) show that aggregated benchmark scores (e.g., BIG-Bench) follow smooth scaling trends across diverse model families, yet individual tasks are more difficult to predict. Ruan et al. (2024) explore observational rather than purely experimental scaling, leveraging over a hundred pretrained LLMs to fit parametric curves without training each from scratch. These lines of research challenge the assumption that scaling laws transfer seamlessly across all model families. Indeed, Choshen et al. (2025) highlight large predictive errors when applying scaling trends derived from one family to a different, structurally distinct family. Overall, while scaling laws are fundamental to LLM research, their reliance on homogeneous or near-homogeneous settings can limit predictive accuracy, especially when addressing newly merged or fine-tuned models for which classical scaling metrics (e.g., total FLOPs or training tokens) are not strictly comparable.
Predicting Downstream Task Performance from Observational Data.
A growing body of work aims to bypass some limitations of traditional scaling laws by using observational or empirical records from existing models. Rather than modeling only cross-entropy loss, these methods directly estimate how well a new or partially evaluated LLM will perform on downstream benchmarks. One common strategy is to reduce the high cost of comprehensive benchmarking by sampling only a small subset of instances. For example, Polo et al. (2024), Kipnis et al. (2025), and Pacchiardi et al. (2024) show that evaluating each new LLM on 100–600 carefully chosen samples can often predict that LLM’s accuracy on thousands of held-out benchmark items with minimal error. Such “tiny benchmarks” mitigate the expense of model-by-model, full-scale evaluations.
Beyond sample-efficient evaluation, another theme is to learn joint representations of both models and tasks. Zhang et al. (2024) combine historical LLM results with model and task descriptors (e.g., parameter size, instruction-tuning data) to improve prediction via matrix factorization or neural collaborative filtering. A similar philosophy underlies the works of Ruan et al. (2024) and Choshen et al. (2025), who incorporate both scaling metrics and observed performance to predict complex “emergent” or “agentic” behaviors of new models. Although these observational approaches can generalize better across different LLM architectures, they seldom exploit fine-grained lineage information—such as which parent checkpoints were merged or how a model was fine-tuned—even though such lineage frequently constrains or correlates performance across tasks.
Routing, Ensemble Methods, and Embedding-Based Prediction.
A complementary direction addresses how to select or route queries among multiple candidate LLMs that each has known strengths. In kNN-based routing (Shnitzer et al., 2024; Hu et al., 2024), for instance, the system uses prompt embeddings or input features to decide which model is most likely to excel on a particular query; Lu et al. (2024) similarly train routing functions based on reward signals. Such approaches implicitly rely on predicting individual models’ performance from prior records, so improved performance estimation methods can feed directly into these routing pipelines. Meanwhile, Ong et al. (2025) propose a factorization approach (combining matrix factorization with label-enhancement techniques) that learns a dynamic “router” capable of choosing between smaller vs. larger models at inference time to optimize cost-performance trade-offs. These routing-oriented works share a common challenge: how to reliably predict an LLM’s performance on a new input. While embedding-based or factor-based methods capture prompt similarity, they again treat each model as a separate entity without leveraging the relationship between parent and child or among merged “soups.”
Motivation for Lineage-Aware Estimation.
Merging or fine-tuning from existing checkpoints is pervasive in LLM development (Wortsman et al., 2022). Two “child” models derived from the same “parent” frequently behave similarly across tasks, and naive averaging of parent checkpoints can sometimes preserve or even enhance performance in certain domains. Nevertheless, conventional scaling laws or standard matrix factorization rarely encode such genealogical constraints. As a result, performance estimates for newly merged or fine-tuned models often rely on expensive direct evaluations or broad assumptions. Our work addresses this gap by building on empirical performance-based frameworks and injecting explicit lineage regularization, thereby offering a more reliable way to predict the behavior of newly derived models without exhaustive re-benchmarking.
3 Method
Our goal is to predict the performance of Large Language Models (LLMs) on specific tasks or benchmark instances by leveraging partial observations and the inherent relationships among models. We first outline existing approaches based purely on observed performance and factors, then propose our lineage-based method.
3.1 Problem Statement
Let $\mathcal{M}$ be the set of LLMs and $\mathcal{X}$ be the set of problems. Define $z_{u,i}\in\{0,1\}$ as the evaluation score for problem $i\in\mathcal{X}$ solved by LLM $u\in\mathcal{M}$. Here, $z_{u,i}=1$ if the LLM correctly solves the problem, and 0 otherwise. For some LLM-problem pairs, their true scores are already known, and using these observed data, our goal is to accurately estimate the scores for unobserved pairs.
3.2 Existing Methods for LLM Performance Prediction
Historical-Performance-Based Methods.
Performance-based prediction methods generally rely on historical performance data. These methods assume that if two models or problems exhibit similar observed performance, their unobserved performances will also be similar. Formally, they define a general function $f(\cdot)$ that estimates performance using latent embeddings or features of models and instances:
$\hat{z}_{u,i}=f(m_{u},x_{i})$
where $m_{u}\in\mathbb{R}^{d}$ and $x_{i}\in\mathbb{R}^{d}$ are embeddings for the model $u$ and instance $i$, respectively. Specific examples include:
First, in the Matrix Factorization (MF) approach, the prediction is computed as
$\hat{z}_{u,i}=\sigma(m_{u}x_{i}^{\top}),$
where $m_{u}$ and $x_{i}$ are latent vectors representing the user and item, respectively, and $\sigma$ is the sigmoid function.
Second, in the Item Response Theory (IRT) model, the prediction is given by
$\hat{z}_{u,i}=\sigma(a_{i}\theta_{u}-b_{i}),$
where $\theta_{u}$ denotes the ability parameter of user $u$, and $a_{i}$ and $b_{i}$ are the discrimination and difficulty parameters of item $i$, respectively.
Finally, the Neural Collaborative Filtering (NCF) model estimates the interaction as
$\hat{z}_{u,i}=\sigma(\text{MLP}(m_{u},x_{i}\mid\theta)),$
where $\text{MLP}(\cdot)$ is a multi-layer perceptron parameterized by $\theta$, applied to the concatenation of $m_{u}$ and $x_{i}$.
To learn these embeddings, methods typically minimize a binary cross-entropy (BCE) loss on observed performance scores, combined with an $L_{2}$ regularization term:
$\mathcal{O}_{\text{BCE}}=-\frac{1}{|\Omega|}\sum_{(u,i)\in\Omega}\left[z_{u,i}\log(\hat{z}_{u,i})+(1-z_{u,i})\log(1-\hat{z}_{u,i})\right]+\lambda_{\text{L2}}(\|M\|_{F}^{2}+\|X\|_{F}^{2})$
where $\Omega$ is the set of observed LLM-instance pairs, $M$ and $X$ are the matrices containing embeddings $m_{u}$ and $x_{i}$ as rows, respectively, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
The methods described above face significant limitations when dealing with new LLMs that have no observed performance data—the well-known cold-start problem in recommendation systems. This issue is particularly acute in LLM development, where developers must often make decisions about newly created models without having the resources to evaluate them extensively.
Factor-Enhanced Method.
Further improvements can be achieved by explicitly incorporating model and task design factors, such as parameter size, dataset scale, and task characteristics, into the prediction (He et al., 2017; Zhang et al., 2024). The enhanced NCF (NCF with factors) includes these additional factors as embedding vectors $e_{v_{i}},e_{v_{j}}$:
$\hat{s}_{ij}=\text{MLP}(p_{i},q_{j},e_{v_{i}},e_{v_{j}}|\theta)$
Specifically, Zhang et al. (2024) considered model-related factors including Model Family (e.g., LLaMA 2, Pythia), Pretraining Dataset Size (tokens), Parameter Size, GPU Time, FLOPs, Context Window Size, Batch Size, Number of Layers, Number of Attention Heads, Key/Value Size, Bottleneck Activation Size, and Carbon Emission. They also included task-related factors such as Ability (e.g., reasoning), Task Family (e.g., ARC), Output Format (e.g., binary classification), and Few-Shot Setting.
Shapley value analysis indicated that among model factors, Pretraining Dataset Size, Model Family, and Batch Size were most critical, while for task factors, Ability was most influential. These results underscore that factors beyond traditional scaling laws significantly influence prediction accuracy.
Notably, in a homogeneous model set derived from the same model family, factors such as Parameter Size, Context Window Size, and Batch Size often remain identical or highly similar across models.
3.3 Proposed: Lineage-Based Prediction
We address the cold-start problem by leveraging an underutilized source of information: the lineage relationships between LLMs. When a new model is derived through fine-tuning or merging existing models, its behavior is likely to be inherited from its ”ancestors” or ”parents.” This intuition forms the foundation of our approach. While factor-based methods (e.g. NCF with factors) attempt to transfer information indirectly by identifying models with similar characteristics (e.g., similar parameter sizes, training data, or architectural features), our lineage-based approach enables direct information transfer along known derivation paths. This distinction is crucial: rather than inferring similarities from potentially noisy or incomplete metadata, we utilize the explicit knowledge of how models are derived from one another, capturing relationships that may not be apparent from factors alone. To leverage lineage information effectively, we propose two distinct approaches. The first approach, Model Lineage Averaging, directly utilizes lineage relationships in their simplest form to make predictions. The second approach, Lineage-Regularized Matrix Factorization (LRMF), integrates lineage constraints into the collaborative filtering framework, allowing for more sophisticated modeling of performance transfer patterns between related models.
Proposed-1: Model Lineage Averaging
Fine-tuned models and merged models inherently inherit the performance characteristics of their base models. We can estimate performance on an instance $i$ for a new or sparsely evaluated LLM $u$ from the results of its lineage-related (neighbor) LLMs:
$\hat{z}_{u,i}\;=\;\frac{1}{|\mathcal{N}_{\mathcal{M}}(u)|}\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{M}}(u)}\;z_{u^{\prime},i}.$
Here, the neighborhood $\mathcal{N}_{\mathcal{M}}(u)$ is defined based on lineage relationships. Specifically, if an LLM $u$ is derived from $v$ through a short lineage path (fine-tuning, merging, etc.), we treat $v$ as a neighbor of $u$. This simple yet effective approach directly addresses the cold-start problem by transferring performance knowledge from established models to their derivatives.
Proposed-2: Lineage-Regularized Matrix Factorization (LRMF).
While derived models fundamentally inherit capabilities from their base models, subtle differences emerge in practice during fine-tuning or merging processes. We account for these nuances by drawing inspiration from Collaborative Filtering approaches. Following the framework of Tag Informed Collaborative Filtering (TagiCoFi) (Zhen et al., 2009), which integrates additional information (e.g., user-generated tags) into the matrix factorization procedure to improve recommendation performance, we propose Lineage-Regularized Matrix Factorization (LRMF). LRMF extends traditional matrix factorization by explicitly encouraging similarity between embeddings of neighboring LLMs and neighboring tasks.
Specifically, let $\mathbf{A}^{(\mathcal{M})}$ and $\mathbf{A}^{(\mathcal{X})}$ be adjacency matrices indicating lineage connections among LLMs and similarity relationships among tasks, respectively. For models, $\mathbf{A}^{(\mathcal{M})}$ directly encodes the lineage relationships: an edge exists between models if one is derived from the other through fine-tuning or merging. For example, if model $v$ is fine-tuned from model $u$, or if model $w$ is created by merging models $u$ and $v$, then $\mathbf{A}^{(\mathcal{M})}_{u,v}=\mathbf{A}^{(\mathcal{M})}_{v,u}=1$ and $\mathbf{A}^{(\mathcal{M})}_{u,w}=\mathbf{A}^{(\mathcal{M})}_{v,w}=1$.
We introduce regularization terms to ensure neighboring embeddings remain close. In particular, for models, the regularization term that enforces similarity between feature vectors of lineage-related models is given by:
$\mathcal{O}^{(\mathcal{M})}_{2}=\frac{1}{2}\sum_{u,v}\mathbf{A}^{(\mathcal{M})}_{u,v}\|\mathbf{m}_{u}-\mathbf{m}_{v}\|^{2}.$
This term penalizes large differences between the embeddings of models that share a lineage connection. Expanding and simplifying this expression yields:
$\mathcal{O}^{(\mathcal{M})}_{2}=\sum_{u}\mathbf{m}_{u}^{\top}\mathbf{m}_{u}\,\mathbf{D}^{(\mathcal{M})}_{u,u}-\sum_{u,v}\mathbf{m}_{u}^{\top}\mathbf{m}_{v}\,\mathbf{A}^{(\mathcal{M})}_{u,v}=\text{Tr}(\mathbf{M}^{\top}\mathbf{L}^{(\mathcal{M})}\mathbf{M}),$
where the diagonal degree matrix $\mathbf{D}^{(\mathcal{M})}$ is defined by $\mathbf{D}^{(\mathcal{M})}_{u,u}=\sum_{v}\mathbf{A}^{(\mathcal{M})}_{u,v}$ and the graph Laplacian is $\mathbf{L}^{(\mathcal{M})}=\mathbf{D}^{(\mathcal{M})}-\mathbf{A}^{(\mathcal{M})}$.
Similarly, for tasks, we construct $\mathbf{A}^{(\mathcal{X})}$ using cosine similarity between task embeddings, retaining the top $k$ similar tasks, and define a corresponding regularization term $\mathcal{O}^{(\mathcal{X})}_{2}=\text{Tr}(\mathbf{X}^{\top}\mathbf{L}^{(\mathcal{X})}\mathbf{X})$.
The overall objective function is then formulated as:
$\mathcal{O}=\mathcal{O}_{\text{BCE}}+\lambda_{\mathcal{X}}\,\mathcal{O}^{(\mathcal{X})}_{2}+\lambda_{\mathcal{M}}\,\mathcal{O}^{(\mathcal{M})}_{2}.$
Here, the regularization parameters $\lambda_{\mathcal{X}}$ and $\lambda_{\mathcal{M}}$ control the influence of instance similarity and model lineage constraints, respectively. A detailed analysis of these hyperparameters and their effects on model performance is provided in Appendix B. To optimize this objective function, we employ a stochastic gradient descent method with adaptive learning rates (Adam optimizer), which efficiently handles the non-convex nature of our formulation while providing fast convergence rates. This approach addresses data sparsity and significantly alleviates cold-start issues for newly derived models.
4 Experiments
We evaluate the effectiveness of our proposed Lineage-Based Prediction methods through extensive empirical experiments using publicly available data from the Hugging Face Open LLM Leaderboard v2 (Fourrier et al., 2024). Our experiments focus on two critical scenarios encountered in real-world LLM development: (1) predicting performance among homogeneous models (e.g., those derived from merging or fine-tuning within the same model family) and (2) predicting performance across more diverse, heterogeneous models (e.g., different architectures or training methods).
4.1 Experimental Setting
We conducted large-scale experiments using evaluations from the Hugging Face Open LLM Leaderboard, focusing on 6 benchmark tasks (BBH (Suzgun et al., 2023), GPQA (Rein et al., 2024), IFEval (Zhou et al., 2023), MATH (Hendrycks et al., 2021b), MMLU-Pro (Wang et al., 2024), MuSR (Sprague et al., 2024)) comprising 39 sub-benchmarks in total. Our dataset includes up to 2,934 publicly available models with partial or complete lineage information and 21,065 instances, though specific subsets were used depending on the experimental condition.
We compared three methods in our experiments:
Baseline: Neural Collaborative Filtering with Factors (NCF with factors).
We used the same network architecture as in previous work (Zhang et al., 2024), consisting of two hidden layers with 128 units each. Due to practical constraints when collecting data from the leaderboard, we limited the model factors to three key attributes: architecture type (50 distinct categories including qwen2forcasualllm, gemma2forcasuallm, etc.), model type (3 categories: Fine-Tuned models, Merged models, and Others including base models), and parameter size. For task factors, we included benchmark identification (6 main benchmarks divided into 39 sub-benchmarks). These factors were encoded as embeddings and concatenated with the model and instance embeddings before being processed by the MLP.
Proposed-1: Model Lineage Averaging.
We constructed lineage relationships by extracting information from Hugging Face model cards. For Fine-Tuned models, we identified parent models from the “base_model:finetune:” tag. For Merged models, we collected parent models from both the “base_model:merge:” tag and the merge_config.json file when available. This approach allowed us to establish direct genealogical connections between derivatives and their parent models, forming the basis for our neighborhood definitions in the lineage averaging process.
Proposed-2: Lineage-Regularized Matrix Factorization.
For our proposed method, we constructed the instance similarity graph using semantic embeddings of the prompts. Specifically, we utilized Snowflake/snowflake-arctic-embed-l-v2.0 to generate high-quality embeddings of the prompt text, and then computed cosine similarities between these embeddings to define edges in the instance graph. For each instance, we retained connections to its top-$k$ most similar instances (with $k=20$ in our experiments). The model similarity graph was constructed using the same lineage information described in the Model Lineage Averaging baseline, where each edge represents a direct derivation relationship between models. The model was implemented in PyTorch and optimized using the Adam optimizer with a learning rate of 0.003. We trained for a maximum of 10,000 epochs with early stopping when no improvement was observed on the validation set for 100 consecutive epochs.
Hyperparameter Tuning
Detailed hyperparameter tuning procedures and their impact on performance are presented in Appendix B. For all methods, we performed grid searches over relevant regularization parameters ($\lambda_{L2}$, $\lambda_{\mathcal{M}}$, $\lambda_{\mathcal{X}}$) and selected the configuration that maximized AUC-ROC on the development set.
Evaluation Metric
For our experiments, we evaluate each method by calculating the correlation coefficient between the estimated performance and the actual performance of LLMs. This metric was chosen because our primary goal is to accurately capture the relative performance ranking among different models rather than precisely matching the absolute performance values. A high correlation coefficient indicates that the method correctly preserves the performance hierarchy among models, which is crucial for identifying promising model variants during development without requiring exhaustive evaluations. We report Pearson correlation coefficients across various benchmarks to assess each method’s predictive capability.
4.2 Experiment 1: Predicting Performance among Homogeneous Models
Scenario Setup
We first examine scenarios involving homogeneous models, i.e., those derived from a common base model through multiple fine-tunings or merges. This setting simulates practical use cases such as determining which model merges or fine-tunings are most promising.
We use 145 models derived from Qwen 2.5-7B, dividing them into 105 models for training, 20 for validation, and 18 for testing. (The detailed configuration of fine-tuned and merged models is listed in the Appendix.) We sample 1,000 instances from the leaderboard benchmarks as the training set, ensuring each model solves all 1,000 instances. This choice follows findings that performance on 100–600 samples reliably predicts overall benchmark performance (Polo et al., 2024; Kipnis et al., 2025).
Table 1: Correlation coefficients between predicted and actual model performance across various benchmarks for a homogeneous set of models derived from Qwen 2.5-7B. Our proposed Lineage-Regularized Matrix Factorization achieves the highest overall correlation (0.719), demonstrating the effectiveness of combining lineage information with collaborative filtering. Note the benchmark-specific patterns: LRMF excels on instruction-following (IFEval) and general reasoning tasks, while Model Lineage Averaging performs better on specialized knowledge (GPQA) and complex reasoning scenarios (MuSR) where performance characteristics tend to be more directly inherited from parent models.
Method all bbh gpqa ifeval math mmlu-pro musr
Lineage-Regularized MF 0.719 0.563 -0.183 0.702 0.324 0.667 0.208
Model Lineage Averaging 0.367 0.412 0.433 0.175 0.175 0.472 0.612
NCF with factors 0.460 0.493 0.310 0.230 0.253 0.506 0.612
Results
Table 1 shows the correlation coefficients between the actual and predicted performance of each LLM across different benchmarks. For overall performance estimation across the entire leaderboard, Lineage-Regularized Matrix Factorization (LRMF) achieved the highest correlation (0.719), followed by NCF with factors (0.460), and Model Lineage Averaging (0.367). These results demonstrate that considering both lineage information and collaborative filtering patterns yields the most accurate predictions.
However, benchmark-specific results reveal notable differences that highlight the distinct characteristics of each method. Model Lineage Averaging performs particularly well on specialized knowledge tasks like GPQA (0.433) and complex reasoning tasks like MuSR (0.612) where it even outperforms the other methods. This suggests that for these particularly challenging benchmarks, fine-tuning or merging tends to produce relatively modest performance changes, making direct ancestor performance a reliable predictor. In contrast, for instruction-following capabilities measured by IFEval, Model Lineage Averaging performs poorly (0.175) while Lineage-Regularized Matrix Factorization excels (0.702). This substantial difference indicates that instruction-following abilities change significantly during fine-tuning or merging, causing direct lineage information alone to be insufficient, whereas the collaborative patterns captured by LRMF provide more relevant signals.
NCF with factors shows intermediate performance across most benchmarks. In this homogeneous experimental setting, where parameter count and model family are identical across all models, the only meaningful factor is the model type (fine-tuned, merged, or other). This limited factor diversity explains why NCF with factors performs relatively well on structural reasoning tasks like BBH (0.493) and knowledge-intensive benchmarks like MMLU-Pro (0.506), but struggles with the more dramatic performance shifts seen in IFEval (0.230).
4.3 Experiment 2: Predicting Performance among Heterogeneous Models
Scenario Setup
To address realistic model merging scenarios, such as Frankenmerge (Wortsman et al., 2022), we examine performance prediction across diverse models with different architectures and parameter sizes. Specifically, we test whether our method remains effective compared to scaling-law-based approaches like NCF with factors, which typically benefit from heterogeneous training sets.
We train using 2,534 models, with 200 models each for validation and testing. As in Experiment 1, we sample 1,000 instances from the leaderboard benchmarks.
Results
As shown in Table 2, Lineage-Regularized Matrix Factorization again achieved the highest overall correlation (0.579), followed by Model Lineage Averaging (0.523) and NCF with factors (0.256). Contrary to our expectations, NCF with factors performed worse compared to the other methods, despite the increased diversity of models potentially providing more informative factors in this heterogeneous setting. This underperformance likely stems from the factor importance bias observed in prior work (Zhang et al., 2024), where SHAP values revealed that certain strong factors like model family dominated the predictions. When such strong explanatory factors exist, the model tends to output similar predictions for all models within the same family without capturing finer distinctions, which can artificially inflate correlation in homogeneous settings but fails in more diverse model ecosystems.
The benchmark-specific trends remain largely consistent with Experiment 1. Lineage-Regularized Matrix Factorization continues to excel on BBH (0.612), MATH (0.626), and MMLU-Pro (0.617), while Model Lineage Averaging demonstrates stronger performance on GPQA (0.480) and MuSR (0.379). This consistent pattern across different experimental settings reinforces our understanding of how different capabilities transfer through model lineage.
Notably, compared to Experiment 1, both Lineage-Regularized Matrix Factorization and NCF with factors show a relative decline in performance compared to Model Lineage Averaging. For instance, on IFEval, the gap between LRMF and Model Lineage Averaging narrowed substantially (0.370 vs. 0.373). This suggests that in heterogeneous scenarios with greater model diversity, the latent feature space must capture significantly more complex relationships. Direct lineage information becomes relatively more reliable as embedding-based methods struggle to represent the nuanced differences across diverse architectures and training approaches in a unified latent space.
Table 2: Correlation coefficients between predicted and actual model performance across various benchmarks for a heterogeneous set of models comprising all models available on the Hugging Face Open LLM Leaderboard. Our LRMF consistently achieves higher prediction accuracy compared to baseline methods, highlighting its effectiveness in leveraging lineage information even among diverse architectures and training strategies.
Method all bbh gpqa ifeval math mmlu-pro musr
Lineage-Regularized MF 0.579 0.612 -0.081 0.370 0.626 0.617 0.058
Model Lineage Averaging 0.523 0.525 0.480 0.373 0.592 0.544 0.379
NCF with factors 0.256 0.272 0.222 0.078 0.211 0.294 0.131
5 Conclusion
Our work demonstrates that lineage relationships—the explicit parent-child connections between derived LLMs—significantly improve performance prediction accuracy. Across experiments with 2,934 models and six major benchmarks, our Lineage-Regularized Matrix Factorization (LRMF) framework achieved up to 7–10 percentage point increases in correlation with actual performance by effectively modeling ancestral connections between models.
Our analysis revealed that while instruction-following abilities undergo substantial changes during fine-tuning or merging, specialized knowledge and complex reasoning capabilities remain more closely tied to parent models. These insights have practical implications for LLM development, allowing researchers to efficiently explore model variants without exhaustive re-benchmarking, thus reducing computational costs and accelerating development cycles.
Limitations include reduced effectiveness for novel architectures without clear lineage and uniform treatment of all lineage connections regardless of modification type. Future work could explore weighted connections and deeper analysis of how specific fine-tuning techniques affect capability inheritance.
As the LLM ecosystem continues expanding through derivatives rather than training from scratch, tracking lineage information becomes increasingly valuable. Our work shows that understanding “where a model comes from” is just as important as knowing “what it is” when predicting LLM performance.
Appendix A Detailed Model Statistics
We present more extensive information about the models used in each experiment. This includes both the partition splits (train / dev / test) and a breakdown by model type: whether a given model checkpoint is (1) a Fine-Tuned (FT) model, derived from a single parent model via continued training on additional or specialized datasets; (2) a Merged model, created through parameter-level averaging or “souping” of multiple parent models; or (3) an Other model, which encompasses base models released without any further fine-tuning or merging, as well as models for which lineage information is incomplete or unavailable.
A.1 Experiment 1: Homogeneous Models (Qwen-2.5-7B Family)
Table 3 lists all 143 models in the Qwen 2.5-7B family used in the homogeneous setting, along with their assigned partition (105 for training, 20 for development, 18 for testing) and their lineage type (Merged / FT / Other). This family traces back to the publicly available Qwen 2.5-7B checkpoint, with each descendant model created via single or multiple merges and/or domain-specific fine-tuning steps.
Lineage Analysis of Qwen 2.5-7B Family
We investigated the lineage distribution within this homogeneous group, discovering that the number of lineage connections roughly follows a logarithmic distribution (Figure 1). Furthermore, we visualized the complete lineage network among the Qwen models (Figure 2), revealing that the model with the most lineage connections is Qwen-2.5-7B-Instruct.
Table 3: Composition of the heterogeneous set used in Experiment 1.
Partition Total FT Models Merged Models Other (including Base)
Train 105 45 (42.9%) 60 (57.1%) 0 (0.0%)
Dev 20 8 (40.0%) 12 (60.0%) 0 (0.0%)
Test 18 9 (50.0%) 9 (50.0%) 0 (0.0%)
Figure 1: Distribution of lineage connections among the Qwen 2.5-7B family. The distribution approximates a logarithmic trend.
Figure 2: Complete lineage network visualization among the Qwen 2.5-7B models. The most densely connected node is Qwen-2.5-7B-Instruct.
A.2 Experiment 2: Heterogeneous Models (All Models)
In the heterogeneous scenario (Experiment2), we used a total of 2,934 models from the Hugging Face Open LLM Leaderboard that included partial or complete lineage metadata. These were split into 2,534 for training, 200 for development, and 200 for testing. Due to the diversity in architectures, sizes, merge strategies, and fine-tuning procedures, we do not list each model name. Instead, Table4 summarizes the approximate proportions of base, merged, and fine-tuned models.
Lineage Analysis of Heterogeneous Models
We also conducted a lineage analysis across the full heterogeneous set, uncovering again a roughly logarithmic distribution in the number of lineage connections (Figure 3). Approximately 60% of these models are either fine-tuned or merged variants, suggesting lineage information is widely applicable for modeling relationships and improving predictions across a broad spectrum of publicly available models.
Table 4: Composition of the heterogeneous set used in Experiment 1.
Partition Total FT Models Merged Models Other (including Base)
Train 2,534 856 (33.8%) 707 (27.9%) 971 (38.3%)
Dev 200 56 (28.0%) 54 (27.0%) 90 (45.0%)
Test 200 54 (27.0%) 66 (33.0%) 80 (40.0%)
Figure 3: Distribution of lineage connections across the heterogeneous model set. A logarithmic distribution is again apparent, indicating many models have relatively few lineage connections, with only a minority having extensive lineages.
Appendix B Hyperparameter Tuning
Our experiments involve performance data collected from six benchmark tasks provided by the Hugging Face Open LLM Leaderboard, totaling approximately 21,000 instances. For each instance, we have binary outcomes indicating whether a given LLM correctly solved the instance or not. The task is thus formulated as predicting these binary instance-level outcomes for model-instance pairs for which no evaluation is available.
Hyperparameter tuning for all methods was performed using the validation set, optimizing for the instance-level AUC-ROC metric, as this metric effectively captures the method’s capability to distinguish between correct and incorrect model predictions. Below, we detail the hyperparameter search procedures and illustrate the effect of varying the number of observed instances ($t$) on prediction quality.
B.1 Lineage-Regularized MF
Lineage-Regularized MF extends the conventional matrix factorization by including additional Laplacian regularization terms:
$\lambda_{\mathcal{M}}\,\mathrm{Tr}\!\left(M^{\top}L^{(\mathcal{M})}M\right)+\lambda_{\mathcal{X}}\,\mathrm{Tr}\!\left(X^{\top}L^{(\mathcal{X})}X\right),$
where $L^{(\mathcal{M})}$ and $L^{(\mathcal{X})}$ are the graph Laplacians for models and instances, respectively, and an $L_{2}$ penalty with coefficient $\lambda_{\mathrm{L2}}$ is also applied. The hyperparameters $\lambda_{\mathrm{L2}},\lambda_{\mathcal{M}},\lambda_{\mathcal{X}}$ were tuned over the range $\{10^{-9},10^{-8},\ldots,10^{0}\}$ using the development set.
Figure 4 illustrates the effect on AUC-ROC when varying $\lambda_{\mathcal{M}}$ (lineage regularization) and $\lambda_{\mathcal{X}}$ (instance similarity regularization), while fixing $\lambda_{\mathrm{L2}}=10^{-5}$.
For the instance similarity regularization parameter $\lambda_{\mathcal{X}}$, we observe that performance remains relatively stable (AUC-ROC $\approx$ 0.60) for values between $10^{-9}$ and $10^{-6}$, then gradually improves to peak at $\lambda_{\mathcal{X}}\approx 10^{-4}$ with an AUC-ROC of approximately 0.62. Beyond this optimal value, performance decreases significantly, dropping to an AUC-ROC of approximately $0.55$ when $\lambda_{\mathcal{X}}\geq 10^{-2}$.
For the lineage regularization parameter $\lambda_{\mathcal{M}}$, we observe remarkably stable performance across all tested values ranging from $10^{-9}$ to $10^{-3}$, with AUC-ROC consistently around $0.62-0.63$. This stability suggests that model lineage information provides robust guidance for performance prediction across a wide range of regularization strengths, with a slight preference for values at the lower end of the spectrum.
These results indicate that while instance similarity regularization requires careful tuning with an optimal value around $10^{-4}$, the lineage-based regularization is considerably more robust to parameter selection. This finding highlights the inherent stability of lineage information as a guiding constraint for performance prediction.
Once the best hyperparameter values are identified on the development set, we retrain Lineage-Regularized MF on the combined training and development sets and then report the final performance on the test set. The same procedure is applied to all baselines to ensure a fair comparison.
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
Figure 4: Development-set AUC-ROC of Lineage-Regularized MF as a function of the lineage regularization parameter $\lambda_{\mathcal{M}}$ and the instance regularization parameter $\lambda_{\mathcal{X}}$. Each plot fixes one parameter while sweeping the other. Optimal performance is achieved with moderate regularization ($\lambda_{\mathcal{M}}\approx 10^{-4}$).
Once the best hyperparameter values are identified on the development set, we retrain Lineage-Regularized MF on the combined training and development sets and then report the final performance on the test set. The same procedure is applied to all baselines to ensure a fair comparison.
Appendix C Effect of the Number of Observed Instances ($t$)
We investigate how varying the number of observed instances per model ($t$) affects predictive performance. By sampling $t\in\{5,10,20,50,100,200,500,1000\}$ instances for each model in the training data, we simulate scenarios ranging from minimal evaluation (resource-constrained) to comprehensive benchmarking.
Figures 5 and 6 present the test-set AUC-ROC across two critical evaluation scenarios:
  • In-distribution prediction: Performance on withheld instance-model pairs where the models appeared in training but certain instances were unobserved (Figure 5(a) and 6(a)).
  • Cold-start prediction: Zero-shot performance on entirely new models unseen during training, representing the challenging cold-start scenario in LLM development (Figure 5(b) and 6(b)).
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
Figure 5: AUC-ROC as a function of observed instances ($t$) per model, evaluated on (a) unobserved instance-model pairs within training models and (b) completely unseen test models for the homogeneous Qwen 2.5-7B family.
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
Figure 6: AUC-ROC as a function of observed instances ($t$) per model, evaluated on (a) unobserved instance-model pairs within training models and (b) completely unseen test models for the heterogeneous model set derived from the Hugging Face Open LLM Leaderboard.
The results align with our main findings in Section 4: Lineage-Regularized MF consistently outperforms baselines, with its advantage becoming more pronounced in the cold-start scenario and with fewer observed instances per model. This confirms our hypothesis that lineage relationships provide valuable signals for predicting the performance of new or sparsely evaluated models.
Impact of Limited Observations
To quantify the robustness of our approach under data-constrained scenarios, we systematically reduced the number of observed instances per model from 1,000 down to as few as 5 instances. As shown in Figure 7, Lineage-Regularized MF maintains impressive prediction accuracy even with drastically reduced observations, demonstrating significant resilience compared to Model Lineage Averaging and NCF with factors.
The performance gap between methods is especially pronounced in the homogeneous setting (Figure 7(b)), where Lineage-Regularized MF achieves with just 50-100 instances per model what baseline methods require 500+ instances to match. This 5-10x reduction in required evaluations translates to substantial computational savings during model development. In the heterogeneous setting (Figure 7(a)), while all methods benefit from the diversity of models, our approach still maintains a consistent advantage across all observation levels.
(a) Heterogeneous Models (All Models)
(b) Homogeous Models (Qwen-2.5-7B)
Figure 7: Correlation between predicted and actual model performance as the number of observed instances per model ($t$) varies. (a) shows results for the heterogeneous set of all models, and (b) for the homogeneous set of models derived from Qwen 2.5-7B. Lineage-Regularized MF maintains consistently high prediction accuracy even with significantly fewer observations, demonstrating robustness compared to baseline methods.
These results have significant practical implications: developers can reliably predict a new model’s performance across multiple benchmarks after evaluating it on just 50-100 carefully selected instances, provided that lineage information is properly incorporated. This dramatically reduces the computational overhead typically associated with comprehensive benchmarking during iterative LLM development.